On four consecutive running plays, Duce Staley of the Philadelphia Eagles gained 23 years, lost 5 yards, gained 15 yards, and lost 10 yards. What was his net yardage change for the series of plays?
Thanks

Answer by checkley71(8403)

23-5+15-10=38-15=23 answer for the total yards gained.